The supply curve for product X is given by QXS = -400 + 10PX .

a. Find the inverse supply curve.


b. How much surplus do producers receive when Qx = 500? When Qx = 1,250?
Answer #1

Answer
a)
the inverse supply curve is
Q=-400+10P
10P=Q+400
P=0.1Q+40
the inverse supply curve is
P=0.1Q+40
-----------
b)
producer surplus =0.5*(Price-Y intercept of supply curve )*Q
Y intercept of the supply curve is 40 from the inverse supply curve

Q=500
P=0.1*500+40=$90
PS=0.5*(90-40)*500
=12500

the producer surplus is $12500
--------
Q=1250
P=0.1*1250+40=165
PS=0.5*(165-40)*1250
=78125

the producer surplus is $78125